# homogeneous_khovanskii.bib

### v1

@@ Abstract.
In 2019, Kaveh and Manon introduced Khovanskii bases as a special ‘Gröbner/like’ generating system of an algebra. We extend their work by considering an arbitrary grading on the algebra and propose a definition for a ‘homogeneous Khovanskii basis’ that respects this grading. We generalize Khovanskii bases further by taking multiple valuations into account (MUVAK bases). We give algorithms in both cases.
@@ 1. Introduction
SAGBI bases – the subalgebra analogue to Gröbner bases for ideals – were introduced by Robbiano and Sweedler in [RS90]. One is given a subalgebra $A\leq K[X_{1},\dots,X_{n}]$ of a polynomial ring over a field $K$ and a monomial ordering $>$. The algorithmic task is to find generators $f_{1},\dots,f_{k}$ of $A$ with
@@
where $\operatorname{LM}_{>}(f)$ denotes the leading monomial of a polynomial $f$ with respect to $>$. SAGBI bases are also called canonical bases in [Stu96]. Kaveh and Manon [KM19] generalized this concept by far with Khovanskii bases. In this setting, $A$ is a finitely generated $K$/̄algebra without zero/divisors and $v:A\setminus\{0\}\to\mathbb{Q}^{r}$ is a discrete valuation on $A$ for some $r\in\mathbb{Z}_{>0}$. The valuation $v$ induces a filtration on $A$ and one considers the associated graded algebra $\operatorname{gr}_{v}(A)$, see below for details. One now wants to find generators of $A$ whose images in $\operatorname{gr}_{v}(A)$ generate $\operatorname{gr}_{v}(A)$ as a $K$/̄algebra.
In this article, we extend the work of Kaveh and Manon and consider a grading on $A$ by an arbitrary finitely generated abelian group $\Delta$. We give a definition of a $\Delta$/̄homogeneous Khovanskii basis, which – as the name suggests – should be a $\Delta$/̄homogeneous generating system of $A$ that fulfils a ‘Khovanskii/like’ property, see section 2 for the technical details. We further bridge the gap between SAGBI bases and Khovanskii bases by allowing an ambient ring $B$ of which $A$ is a subalgebra. In case of SAGBI bases, $B$ would be a polynomial ring; for a Khovanskii basis as in [KM19], we have $A=B$.
We further introduce a ‘Multiple Valuations Analogue to Khovanskii bases’ (MUVAK bases), where a Khovanskii/like property is required to be fulfilled for several valuations simultaneously. Our construction is motivated by the computation of certain Cox rings. In the past decade, there has been quite some interest in Cox rings of minimal models of quotient singularities [Don16, DW17, Yam18, Gra19, Sch23] and their algorithmic construction contributes to the aim of having an ‘algorithmic minimal model programme’, see [Laz24]. In [Yam18], Yamagishi gives an algorithm to compute generators of such a Cox ring and, in the language of this article, the algorithm in fact computes a MUVAK basis.
We describe our setting and give the definition of a homogeneous Khovanskii basis in section 2. We further present a ‘homogeneous version’ of the subduction algorithm (Algorithm 1), that is, an algorithm to reduce elements by generators of a subalgebra. As in [KM19], this algorithm does in general not terminate. However, we are able to prove its termination for every input under the condition that any $\Delta$/̄graded component of $A$ is of finite dimension (Proposition 2.13), which notably includes the setting of a polynomial ring with the standard grading. We further give an algorithm to compute a homogeneous Khovanskii basis, see Algorithm 2. The algorithm is similar to the one given in [KM19] and relies in the same way on the equality of two ideals. We prove the corresponding theorem (Theorem 2.19) only under the assumption that the subduction algorithm (provably) terminates.
@@
We finally discuss our initial motivation for this article, namely the computation of Cox rings of minimal models of quotient singularities in section 5. We recall a theorem by Yamagishi [Yam18] and Grab [Gra19] that characterizes a generating system for such a Cox ring by a MUVAK bases. In [Yam18], gives an algorithm to compute such generators. We implemented the algorithms in this paper in the computer algebra system OSCAR [Osc, DEFHJ24] and compared these approaches with our implementation of Yamagishi’s algorithm stemming from [Sch23]. Our algorithms appear to be an improvement of Yamagishi’s algorithm, see the timings in Table 1.
In the appendix, we describe an efficient method to compute the homogenization of an ideal with respect to non/negative weights by a generalization of ‘Bayer’s method’. This is relevant for a performant implementation of both Algorithm 3 and Yamagishi’s algorithm.
@@ Acknowledgements
I thank Tommy Hofmann for helpful comments on an early version of this article. This work was supported by the SFB/TRR 195 ‘Symbolic Tools in Mathematics and their Application’ of the German Research Foundation (DFG).
@@ 2.1. Setting
Throughout, let $K$ be a field and let $A\leq B$ be finitely generated $K$/̄algebras and domains. Let $(\Gamma,+,\succeq)$ be a linearly ordered group, that is, an abelian group $(\Gamma,+)$ equipped with a total ordering $\succeq$ such that $\gamma_{1}\succeq\gamma_{2}$ implies $\gamma_{1}+\gamma_{3}\succeq\gamma_{2}+\gamma_{3}$ for all $\gamma_{1},\gamma_{2},\gamma_{3}\in\Gamma$.
@@
A function $v:B\setminus\{0\}\to\Gamma$ is a valuation over $K$ if it satisfies the following axioms:
@@
    For all $0\neq f,g\in B$ with $0\neq f+g$ we have $v(f+g)\succeq\min\{v(f),v(g)\}$.
@@
    For all $0\neq f\in B$ and $0\neq\lambda\in K$ we have $v(\lambda f)=v(f)$.
In the following, let $v:B\setminus\{0\}\to\Gamma$ be a valuation over $K$. The valuation $v$ gives a $\Gamma$/̄filtration $\mathcal{F}_{v}=(F_{v\succeq\gamma})_{\gamma\in\Gamma}$ on $B$ in the following way. For $\gamma\in\Gamma$, we set
@@
    Let $f\in A\setminus\{0\}$ be $\Delta$/̄homogeneous of degree $\deg_{\Delta}(f)=\delta$. We define
@@
[…]
@@
Let $\mathcal{B}\subseteq A$ be a set of $\Delta$/̄homogeneous generators of $A$ as a $K$/̄algebra. We call $\mathcal{B}$ a $\Delta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$, if the set $\{\operatorname{in}_{v}^{\Delta}(f)\mid f\in\mathcal{B}\}$ generates $\operatorname{in}_{v}^{\Delta}(A)$ as a $K$/̄algebra.
@@
$\gamma_{1}\succeq\gamma_{2}:\Longleftrightarrow X^{\gamma_{1}}\leq X^{\gamma_{2}}$
for $\gamma_{1},\gamma_{2}\in\mathbb{Z}^{n}$, where by $X^{\gamma_{1}}$ we mean the monomial with exponents $\gamma_{1}$. This gives a valuation $v:B\setminus\{0\}\to\mathbb{Z}^{n}$ over $K$ by putting $v(f)=\gamma$ for $f\in B\setminus\{0\}$ with leading monomial $\operatorname{LM}_{\geq}(f)=X^{\gamma}$. Notice that by construction of $\succeq$, the valuation of $f$ is the smallest exponent among the monomials of $f$ with respect to $\succeq$. In this way, the (ascending) filtration one associates to $\geq$, see [KR05, Example 6.5.4], is equivalent to the (descending) filtration we associate to $v$. With this construction, we have $\operatorname{gr}_{v}(B)=B$ and $\operatorname{in}_{v}(f)=\operatorname{LM}_{\geq}(f)$ for all $f\in B$.
@@ 2.2. The homogeneous subduction algorithm
An essential tool for the computation of SAGBI or Khovanskii bases is the so/called subduction algorithm [RS90, 1.5], [Stu96, Algorithm 11.1], [KM19, Algorithm 2.11]. This algorithm is an analogue of division with remainder in a polynomial ring and allows one to write an element of the algebra as a polynomial in the given basis. We now give a $\Delta$/̄homogeneous version of this algorithm.
@@
Notation 2.8.
For a polynomial $h=\sum_{a\in\mathbb{Z}_{\geq 0}^{n}}\lambda_{a}X_{1}^{a_{1}}\cdots X_{n}^{a_{n}}\in K[X_{1},\dots,X_{n}]$, we denote by $\operatorname{Supp}(h)=\{\lambda_{a}X_{1}^{a_{1}}\cdots X_{n}^{a_{n}}\mid\lambda_{a}\neq 0\}$ the support of $h$, that is, the set of non/̄zero terms.
We present the homogeneous subduction algorithm in Algorithm 1. For full generality, we state the algorithm with a possibly infinite set $\mathcal{B}$. In the algorithm, we work with subrings of the polynomial ring $K[X_{f}\mid f\in\mathcal{B}]$ with variables labelled by the elements of $\mathcal{B}$, so $X_{f_{i}}$ denotes a variable corresponding to $f_{i}\in\mathcal{B}$.
Input : A $\Delta$/̄homogeneous element $f\in A$ and a set $\mathcal{B}\subseteq A$ of $\Delta$/̄homogeneous elements
Output : A finite subset $\{f_{1},\dots,f_{k}\}\subseteq\mathcal{B}$, a polynomial $h\in K[X_{1},\dots,X_{k}]$ and an element $r\in A$ which fulfil all of the following conditions:
@@
    either $r=0$ or $r$ is $\Delta$/̄homogeneous with $\deg_{\Delta}(r)=\deg_{\Delta}(f)$,
@@
    either $r=0$ or $\operatorname{in}_{v}^{\Delta}(r)\notin\operatorname{in}_{v}^{\Delta}(\mathcal{B})$.
1 $\mathcal{B}^{\prime}\coloneqq\emptyset$
@@ 3 $r\coloneqq f$
4 while $r\neq 0$ and $\operatorname{in}_{v}^{\Delta}(r)\in\operatorname{in}_{v}^{\Delta}(\mathcal{B})$ do
5      Find a finite subset $\{f_{1},\dots,f_{k}\}\subseteq\mathcal{B}$ and a polynomial $h^{\prime}\in K[X_{f_{1}},\dots,X_{f_{k}}]$ with $h^{\prime}(\operatorname{in}_{v}^{\Delta}(f_{1}),\dots,\operatorname{in}_{v}^{\Delta}(f_{k}))=\operatorname{in}_{v}^{\Delta}(r)$, $v(t(f_{1},\dots,f_{k}))\succeq v(r)$ and $\deg_{\Delta}(t(f_{1},\dots,f_{k}))=\deg_{\Delta}(r)$ for every $t\in\operatorname{Supp}(h^{\prime})$
6      $\mathcal{B}^{\prime}\coloneqq\mathcal{B}^{\prime}\cup\{f_{1},\dots,f_{k}\}$
@@ 8      $h\coloneqq h+h^{\prime}$
return $\mathcal{B}^{\prime}$, $h$, $r$
@@
Note that the conditions on the terms of $h^{\prime}$ in line 5 of Algorithm 1 can be easily achieved. By assumption, $\operatorname{in}_{v}^{\Delta}(r)\in\operatorname{in}_{v}^{\Delta}(\mathcal{B})$, so there is a finite set $\{f_{1},\dots,f_{k}\}\subseteq\mathcal{B}$ and a polynomial $h^{\prime}\in K[X_{1},\dots,X_{k}]$ with
@@
We endow the polynomial ring $K[X_{1},\dots,X_{k}]$ with a $\Gamma$- and a $\Delta$/̄grading by putting $\deg_{\Gamma}(X_{i})\coloneqq v(f_{i})$ and $\deg_{\Delta}(X_{i})\coloneqq\deg_{\Delta}(f_{i})$. We can now remove all terms $t\in\operatorname{Supp}(h^{\prime})$ from $h^{\prime}$ with $\deg_{\Gamma}(t)\neq v(r)$ or $\deg_{\Delta}(t)\neq\deg_{\Delta}(r)$. This does not change $h^{\prime}(\operatorname{in}_{v}^{\Delta}(f_{1}),\dots,\operatorname{in}_{v}^{\Delta}(f_{k}))$ by $\Gamma$- and $\Delta$/̄homogeneity of $\operatorname{in}_{v}^{\Delta}(r)$. In other words, $h^{\prime}$ is a homogeneous preimage of $\operatorname{in}_{v}^{\Delta}(r)$ under the $(\Gamma,\Delta)$/̄graded morphism given by $X_{i}\mapsto\operatorname{in}_{v}^{\Delta}(f_{i})$.
@@
Assume the algorithm terminates for a given input $f$ and $\mathcal{B}$ with a triple $\mathcal{B}^{\prime}$, $h$ and $r$. We need to show that $\mathcal{B}^{\prime}$, $h$ and $r$ fulfil the conditions (a) to (e) in Algorithm 1.
If the algorithm terminates, condition (e) is fulfilled by construction of the while/̄loop. Further, condition (d) follows from (a) and (b): indeed, (b) implies that $h(f_{1},\dots,f_{k})$ is $\Delta$/̄homogeneous of degree $\deg_{\Delta}(f)$ whenever $h\neq 0$, hence so is $r=f-h(f_{1},\dots,f_{k})$ by (a).
The conditions (a) to (c) are certainly fulfilled before the start of the while/̄loop. We prove that they are maintained during the run of the loop, that is, that after each iteration of the loop (a) to (c) hold. Let $\mathcal{B}^{\prime}$, $h$ and $r$ be the values at the beginning of a given iteration of the loop and let $h^{\prime}$ be the polynomial chosen during this iteration. Condition (a) is maintained by construction. By (c), we have $v(r)\succeq v(f)$ at the beginning of the loop, hence $v(t(f_{1},\dots,f_{k}))\succeq v(f)$ for every $t\in\operatorname{Supp}(h^{\prime})$ and this gives (b) (see Remark 2.9 for the feasibility of line 5 of the algorithm).
@@
In general, there is no guarantee that Algorithm 1 terminates, see Example 4.1. We say that Algorithm 1 terminates for $A$, if the algorithm terminates for every $\Delta$/̄homogeneous $f\in A$ and every set $\mathcal{B}\subseteq A$ of $\Delta$/̄homogeneous elements after finitely many steps. We now give some conditions that ensure the termination.
@@
Assume that the set $S(A,v)\coloneqq\{v(f)\mid 0\neq f\in A\}$ is maximum well/ordered, that is, every subset of $S(A,v)$ has a maximal element with respect to $\succeq$. As in [KM19, Proposition 2.13], we see that Algorithm 1 terminates for $A$ in this case as the valuation of $r$ properly increases in every iteration.
If $B$ is a polynomial ring and $v$ is induced by a monomial ordering $>$ as in Remark 2.7, then the maximum well/ordered property corresponds to $>$ being a (minimum) well/ordering. In other words, $>$ is a global ordering (in the terminology of [GP08]) and termination is guaranteed in the setting of SAGBI bases.
@@
Let $0\neq V\leq A$ be a finite/dimensional $K$/̄vector subspace of $A$. Then there is $\gamma\in\Gamma$ such that $v(a)\preceq\gamma$ for all $a\in V\setminus\{0\}$.
@@
For $\dim(V)=1$, there is nothing to prove. Let $\dim(V)=k$ and let $a_{1},\dots,a_{k}\in V$ be a basis. Write $V^{\prime}\coloneqq\langle a_{1},\dots,a_{k-1}\rangle_{K}$ for the subspace of dimension $k-1$. By induction, there exists $\gamma\in\Gamma$ such that $v(a)\preceq\gamma$ for all $a\in V^{\prime}\setminus\{0\}$. If also $v(a)\preceq\gamma$ for all $a\in V\setminus\{0\}$, we are done, so assume there are $\lambda_{1},\dots,\lambda_{k}\in K$, not all 0, such that for $a\coloneqq\sum_{i=1}^{k}\lambda_{i}a_{i}$ we have $v(a)\succ\gamma$. In particular, $a\notin V^{\prime}$, so $\lambda_{k}\neq 0$.

### v2

# Homogeneous Khovanskii bases and MUVAK bases

@@ Abstract.
In 2019, Kaveh and Manon introduced Khovanskii bases as a special ‘Gröbner-like’ generating system of an algebra. We extend their work by considering an arbitrary grading on the algebra and propose a definition for a ‘homogeneous Khovanskii basis’ that respects this grading. We generalize Khovanskii bases further by taking multiple valuations into account (MUVAK bases). We give algorithms in both cases.
@@ 1. Introduction
SAGBI bases – the subalgebra analogue to Gröbner bases for ideals – were introduced by Robbiano and Sweedler in [RS90]. One is given a subalgebra $A\leq K[X_{1},\dots,X_{n}]$ of a polynomial ring over a field $K$ and a monomial ordering >>. The algorithmic task is to find generators $f_{1},\dots,f_{k}$ of $A$ with
@@
where $\operatorname{LM}_{>}(f)$ denotes the leading monomial of a polynomial $f$ with respect to >>. SAGBI bases are also called canonical bases in [Stu96]. Kaveh and Manon [KM19] generalized this concept by far with Khovanskii bases. In this setting, $A$ is a finitely generated $K$-algebra without zero-divisors and $v:A\setminus\{0\}\to\mathbb{Q}^{r}$ is a discrete valuation on $A$ for some $r\in\mathbb{Z}_{>0}$. The valuation $v$ induces a filtration on $A$ and one considers the associated graded algebra $\operatorname{gr}_{v}(A)$, see below for details. One now wants to find generators of $A$ whose images in $\operatorname{gr}_{v}(A)$ generate $\operatorname{gr}_{v}(A)$ as a $K$-algebra.
In this article, we extend the work of Kaveh and Manon and consider a grading on $A$ by an arbitrary finitely generated abelian group $\Delta$. We give a definition of a $\Delta$-homogeneous Khovanskii basis, which – as the name suggests – should be a $\Delta$-homogeneous generating system of $A$ that fulfils a ‘Khovanskii-like’ property, see section 2 for the technical details. We further bridge the gap between SAGBI bases and Khovanskii bases by allowing an ambient ring $B$ of which $A$ is a subalgebra. In case of SAGBI bases, $B$ would be a polynomial ring; for a Khovanskii basis as in [KM19], we have $A=B$.
We further introduce a ‘Multiple Valuations Analogue to Khovanskii bases’ (MUVAK bases), where a Khovanskii-like property is required to be fulfilled for several valuations simultaneously. Our construction is motivated by the computation of certain Cox rings. In the past decade, there has been quite some interest in Cox rings of minimal models of quotient singularities [Don16, DW17, Yam18, Gra19, Sch23] and their algorithmic construction contributes to the aim of having an ‘algorithmic minimal model programme’, see [Laz24]. In [Yam18], Yamagishi gives an algorithm to compute generators of such a Cox ring and, in the language of this article, the algorithm in fact computes a MUVAK basis.
We describe our setting and give the definition of a homogeneous Khovanskii basis in section 2. We further present a ‘homogeneous version’ of the subduction algorithm (Algorithm 1), that is, an algorithm to reduce elements by generators of a subalgebra. As in [KM19], this algorithm does in general not terminate. However, we are able to prove its termination for every input under the condition that any $\Delta$-graded component of $A$ is of finite dimension (Proposition 2.15), which notably includes the setting of a polynomial ring with the standard grading. We further give an algorithm to compute a homogeneous Khovanskii basis, see Algorithm 2. The algorithm is similar to the one given in [KM19] and relies in the same way on the equality of two ideals. We prove the corresponding theorem (Theorem 2.21) only under the assumption that the subduction algorithm (provably) terminates.
@@
We finally discuss our initial motivation for this article, namely the computation of Cox rings of minimal models of quotient singularities in section 5. We recall a theorem by Yamagishi [Yam18] and Grab [Gra19] that characterizes a generating system for such a Cox ring by a MUVAK basis. In [Yam18], Yamagishi gives an algorithm to compute such generators. We implemented the algorithms in this paper in the computer algebra system OSCAR [Osc25, Dec+25] and compared these approaches with our implementation of Yamagishi’s algorithm stemming from [Sch23]. Our algorithms appear to be an improvement of Yamagishi’s algorithm, see the timings in Table 1. The code used for these timings is freely available online at
https://gitlab.com/math5724907/homogeneouskhovanskii.
In the appendix, we describe an efficient method to compute the homogenization of an ideal with respect to non-negative weights by a generalization of ‘Bayer’s method’. This is relevant for a performant implementation of both Algorithm 3 and Yamagishi’s algorithm.
@@ Acknowledgements
I thank Tommy Hofmann for helpful comments on an early version of this article. This work was supported by the SFB-TRR 195 ‘Symbolic Tools in Mathematics and their Application’ of the German Research Foundation (DFG).
@@ 2.1. Setting
Throughout, let $K$ be a field and let $A\leq B$ be finitely generated $K$-algebras and domains. Let $(\Gamma,+,\succeq)$ be a linearly ordered group, that is, an abelian group $(\Gamma,+)$ equipped with a total ordering $\succeq$ such that $\gamma_{1}\succeq\gamma_{2}$ implies $\gamma_{1}+\gamma_{3}\succeq\gamma_{2}+\gamma_{3}$ for all $\gamma_{1},\gamma_{2},\gamma_{3}\in\Gamma$.
@@
A function $v:B\setminus\{0\}\to\Gamma$ is a valuation over $K$ if $v$ satisfies the following axioms:
@@
    For all $0\neq f,g\in B$ with $0\neq f+g$, we have $v(f+g)\succeq\min\{v(f),v(g)\}$.
@@
    For all $0\neq f\in B$ and $0\neq\lambda\in K$, we have $v(\lambda f)=v(f)$.
In the following, let $v:B\setminus\{0\}\to\Gamma$ be a valuation over $K$. The valuation $v$ gives a $\Gamma$-filtration $\mathcal{F}_{v}=(F_{v\succeq\gamma})_{\gamma\in\Gamma}$ on $B$ in the following way. For $\gamma\in\Gamma$, we set
@@
    Let $f\in A\setminus\{0\}$ be $\Delta$-homogeneous of degree $\deg_{\Delta}(f)=\delta\in\Delta$. We define
@@
For a subset $S\subseteq A$, we denote by $\operatorname{in}_{v}^{\Delta}(S)\leq\operatorname{gr}_{v}(B)\otimes_{K}K\Delta$ the subalgebra generated by the set $\{\operatorname{in}_{v}^{\Delta}(f)\mid f\in S\text{ is $\Delta$\penalty 10000\hbox{-}\penalty 10000\thinspace\ignorespaces homogeneous}\}$.
@@
Let $\mathcal{G}\subseteq A$ be a set of $\Delta$-homogeneous generators of $A$ as a $K$-algebra. We call $\mathcal{G}$ a $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$, if the set $\{\operatorname{in}_{v}^{\Delta}(f)\mid f\in\mathcal{G}\}$ generates $\operatorname{in}_{v}^{\Delta}(A)$ as a $K$-algebra.
@@
Example 2.5.
We use the following running example to illustrate the results in this section. To help with navigation between the examples, we note that this example is continued in 2.12 and 2.23. Let $K$ be a field of characteristic different from 2, $B=K[x,y]$ and $v:B\setminus\{0\}\to\mathbb{Z}$ the order of divisibility by $x$ with the natural ordering on $\mathbb{Z}$. We have
$\operatorname{gr}_{v}(B)=K[y]\oplus xK[y]\oplus x^{2}K[y]\oplus\cdots=B.$
Let $f_{1}\coloneqq x+y,f_{2}\coloneqq x-y\in B$ and set $A=K[f_{1},f_{2}]$. Clearly, $A=B$. We endow $A$ with a grading by $\Delta\coloneqq\mathbb{Z}/2\mathbb{Z}$ via $\deg_{\Delta}(f_{1})=\bar{0}$ and $\deg_{\Delta}(f_{2})=\bar{1}$. We have $v(f_{1})=v(f_{2})=0$ and compute the initial forms:
$\begin{array}[]{c|c}&\operatorname{in}_{v}^{\Delta}\\ \hline\cr f_{1}&y\otimes\bar{0}\\ f_{2}&-y\otimes\bar{1}\end{array}$
We claim that $\{f_{1},f_{2}\}$ is not a $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$. Indeed, $xy=\frac{1}{4}(f_{1}^{2}-f_{2}^{2})\in A$ is homogeneous of degree $\deg_{\Delta}(xy)=\bar{0}$ and hence $\operatorname{in}_{v}^{\Delta}(xy)=xy\otimes\bar{0}\in\operatorname{in}_{v}^{\Delta}(A)$. However, the initial forms $\operatorname{in}_{v}^{\Delta}(f_{i})$ do not generate $xy\otimes\bar{0}$.
@@
$\gamma_{1}\succeq\gamma_{2}\Longleftrightarrow X^{\gamma_{1}}\leq X^{\gamma_{2}}$
for $\gamma_{1},\gamma_{2}\in\mathbb{Z}^{n}$, where by $X^{\gamma_{i}}$ we mean the monomial with exponents $\gamma_{i}$. This gives a valuation $v:B\setminus\{0\}\to\mathbb{Z}^{n}$ over $K$ by putting $v(f)=\gamma$ for $f\in B\setminus\{0\}$ with leading monomial $\operatorname{LM}_{\geq}(f)=X^{\gamma}$. Notice that by construction of $\succeq$, the valuation of $f$ is the smallest exponent among the monomials of $f$ with respect to $\succeq$. In this way, the (ascending) filtration one associates to $\geq$, see [KR05, Example 6.5.4], is equivalent to the (descending) filtration we associate to $v$. With this construction, we have $\operatorname{gr}_{v}(B)=B$ and $\operatorname{in}_{v}(f)=\operatorname{LM}_{\geq}(f)$ for all $f\in B$.
@@ 2.2. The homogeneous subduction algorithm
An essential tool for the computation of SAGBI or Khovanskii bases is the so-called subduction algorithm [RS90, 1.5], [Stu96, Algorithm 11.1], [KM19, Algorithm 2.11]. This algorithm is an analogue of division with remainder in a polynomial ring and allows one to write an element of the algebra as a polynomial in the given basis. We now give a $\Delta$-homogeneous version of this algorithm.
@@
Notation 2.9.
For a polynomial $h=\sum_{a\in\mathbb{Z}_{\geq 0}^{n}}\lambda_{a}X_{1}^{a_{1}}\cdots X_{n}^{a_{n}}\in K[X_{1},\dots,X_{n}]$, we denote by $\operatorname{Supp}(h)=\{\lambda_{a}X_{1}^{a_{1}}\cdots X_{n}^{a_{n}}\mid\lambda_{a}\neq 0\}$ the support of $h$, that is, the set of non-zero terms.
We present the homogeneous subduction algorithm in Algorithm 1. For full generality, we state the algorithm with a possibly infinite set $\mathcal{G}$. In the algorithm, we work with subrings of the polynomial ring $K[X_{f}\mid f\in\mathcal{G}]$ with variables labelled by the elements of $\mathcal{G}$, so $X_{f_{i}}$ denotes a variable corresponding to $f_{i}\in\mathcal{G}$.
Input : A $\Delta$-homogeneous element $f\in A$ and a set $\mathcal{G}\subseteq A$ of $\Delta$-homogeneous elements
Output : A finite subset $\{f_{1},\dots,f_{k}\}\subseteq\mathcal{G}$, a polynomial $h\in K[X_{1},\dots,X_{k}]$ and an element $r\in A$, which fulfil all of the following conditions:
@@
    either $r=0$ or $r$ is $\Delta$-homogeneous with $\deg_{\Delta}(r)=\deg_{\Delta}(f)$,
@@
    either $r=0$ or $\operatorname{in}_{v}^{\Delta}(r)\notin\operatorname{in}_{v}^{\Delta}(\mathcal{G})$.
1
[…]
@@ 4 $r\coloneqq f$
5 while $r\neq 0$ and $\operatorname{in}_{v}^{\Delta}(r)\in\operatorname{in}_{v}^{\Delta}(\mathcal{G})$ do
6   Find a finite subset $\{f_{1},\dots,f_{k}\}\subseteq\mathcal{G}$ and a polynomial $h^{\prime}\in K[X_{f_{1}},\dots,X_{f_{k}}]$ with $h^{\prime}(\operatorname{in}_{v}^{\Delta}(f_{1}),\dots,\operatorname{in}_{v}^{\Delta}(f_{k}))=\operatorname{in}_{v}^{\Delta}(r)$, $v(t(f_{1},\dots,f_{k}))\succeq v(r)$ and $\deg_{\Delta}(t(f_{1},\dots,f_{k}))=\deg_{\Delta}(r)$ for every $t\in\operatorname{Supp}(h^{\prime})$
7   $\mathcal{G}^{\prime}\coloneqq\mathcal{G}^{\prime}\cup\{f_{1},\dots,f_{k}\}$
@@ 9   $h\coloneqq h+h^{\prime}$
return $\mathcal{G}^{\prime}$, $h$, $r$
@@
Note that the conditions on the terms of $h^{\prime}$ in line 5 of Algorithm 1 can be easily achieved. By assumption, $\operatorname{in}_{v}^{\Delta}(r)\in\operatorname{in}_{v}^{\Delta}(\mathcal{G})$, so there is a finite set $\{f_{1},\dots,f_{k}\}\subseteq\mathcal{G}$ and a polynomial $h^{\prime}\in K[X_{1},\dots,X_{k}]$ with
@@
We endow the polynomial ring $K[X_{1},\dots,X_{k}]$ with a $\Gamma$- and a $\Delta$-grading by putting $\deg_{\Gamma}(X_{i})\coloneqq v(f_{i})$ and $\deg_{\Delta}(X_{i})\coloneqq\deg_{\Delta}(f_{i})$. We can now remove all terms $t\in\operatorname{Supp}(h^{\prime})$ from $h^{\prime}$ with $\deg_{\Gamma}(t)\neq v(r)$ or $\deg_{\Delta}(t)\neq\deg_{\Delta}(r)$. This does not change $h^{\prime}(\operatorname{in}_{v}^{\Delta}(f_{1}),\dots,\operatorname{in}_{v}^{\Delta}(f_{k}))$ by $\Gamma$- and $\Delta$-homogeneity of $\operatorname{in}_{v}^{\Delta}(r)$. In other words, $h^{\prime}$ is a homogeneous preimage of $\operatorname{in}_{v}^{\Delta}(r)$ under the $(\Gamma,\Delta)$-graded morphism given by $X_{i}\mapsto\operatorname{in}_{v}^{\Delta}(f_{i})$.
@@
Assume the algorithm terminates for a given input $f$ and $\mathcal{G}$ with a triple $\mathcal{G}^{\prime}$, $h$ and $r$. We need to show that $\mathcal{G}^{\prime}$, $h$ and $r$ fulfil the conditions (a) to (e) in Algorithm 1.
If the algorithm terminates, condition (e) is fulfilled by construction of the while-loop. Further, condition (d) follows from (a) and (b): indeed, (b) implies that $h(f_{1},\dots,f_{k})$ is $\Delta$-homogeneous of degree $\deg_{\Delta}(f)$ whenever $h\neq 0$, hence so is $r=f-h(f_{1},\dots,f_{k})$ by (a).
The conditions (a) to (c) are certainly fulfilled before the start of the while-loop. We prove that they are maintained during the run of the loop, that is, that after each iteration of the loop (a) to (c) hold. Let $\mathcal{G}^{\prime}$, $h$ and $r$ be the values at the beginning of a given iteration of the loop and let $h^{\prime}$ be the polynomial chosen during this iteration. Condition (a) is maintained by construction. By (c), we have $v(r)\succeq v(f)$ at the beginning of the loop, hence $v(t(f_{1},\dots,f_{k}))\succeq v(f)$ for every $t\in\operatorname{Supp}(h^{\prime})$ and this gives (b) (see Remark 2.10 for the feasibility of line 5 of the algorithm).
@@
Example 2.12.
This example is part of the series also containing 2.5 and 2.23. To illustrate Algorithm 1, we run it on a small example. Let
$f=\frac{1}{4}(f_{1}^{2}-f_{2}^{2})+f_{1}=xy+x+y\in A$
and $\mathcal{G}=\{f_{1},f_{2}\}$. We see that $\deg_{\Delta}(f)=\bar{0}\in\mathbb{Z}/2\mathbb{Z}$. Then $\operatorname{in}_{v}^{\Delta}(f)=y\otimes\bar{0}=\operatorname{in}_{v}^{\Delta}(f_{1})$, so the first iteration of the subduction algorithm yields $\mathcal{G}^{\prime}=\{f_{1}\}$, $h=X_{1}$ and $r=f-f_{1}=xy$. Now $\operatorname{in}_{v}^{\Delta}(f)=xy\otimes\bar{0}\notin\operatorname{in}_{v}^{\Delta}(\mathcal{G})$ and the algorithm terminates with non-trivial remainder $r$.
In general, there is no guarantee that Algorithm 1 terminates, see Example 4.1. We say that Algorithm 1 terminates for $A$, if the algorithm terminates for every $\Delta$-homogeneous $f\in A$ and every set $\mathcal{G}\subseteq A$ of $\Delta$-homogeneous elements after finitely many steps. We now give some conditions that ensure the termination.
@@
Assume that the set $S(A,v)\coloneqq\{v(f)\mid 0\neq f\in A\}$ is maximum well-ordered, that is, every subset of $S(A,v)$ has a maximal element with respect to $\succeq$. As in [KM19, Proposition 2.13], we see that Algorithm 1 terminates for AA in this case as the valuation of $r$ properly increases in every iteration.
If $B$ is a polynomial ring and $v$ is induced by a monomial ordering >> as in Remark 2.8, then the maximum well-ordered property corresponds to >> being a (minimum) well-ordering. In other words, >> is a global ordering (in the terminology of [GP08]) and termination is guaranteed in the setting of SAGBI bases.
@@
Let $0\neq V\leq A$ be a finite-dimensional $K$-vector subspace of $A$. Then there is $\gamma\in\Gamma$ such that $v(a)\preceq\gamma$ for all $a\in V\setminus\{0\}$.
@@
For $\dim(V)=1$, there is nothing to prove. Let $\dim(V)=k$ and let $a_{1},\dots,a_{k}\in V$ be a basis. Write $V^{\prime}\coloneqq\langle a_{1},\dots,a_{k-1}\rangle_{K}$ for the subspace of dimension k1k-1. By induction, there exists $\gamma\in\Gamma$ such that $v(a)\preceq\gamma$ for all $a\in V^{\prime}\setminus\{0\}$. If also $v(a)\preceq\gamma$ for all $a\in V\setminus\{0\}$, we are done, so assume there are $\lambda_{1},\dots,\lambda_{k}\in K$, not all 0, such that for $a\coloneqq\sum_{i=1}^{k}\lambda_{i}a_{i}$ we have $v(a)\succ\gamma$. In particular, $a\notin V^{\prime}$, so $\lambda_{k}\neq 0$.